# Evaluation of ChatGPT Usability as A Code Generation Tool

### v2

@@ Abstract
With the rapid advance of machine learning (ML) technology, large language models (LLMs) are increasingly explored as an intelligent tool to generate program code from natural language specifications. However, existing evaluations of LLMs have focused on their capabilities in comparison with humans. It is desirable to evaluate their usability when deciding on whether to use a LLM in software production. This paper proposes a user centric method. It includes metadata in the test cases of a benchmark to describe their usages, conducts testing in a multi-attempt process that mimic the uses of LLMs, measures LLM generated solutions on a set of quality attributes that reflect usability, and evaluates the performance based on user experiences in the uses of LLMs as a tool. The paper reports an application of the method in the evaluation of ChatGPT usability as a code generation tool for the R programming language. Our experiments demonstrated that ChatGPT is highly useful for generating R program code although it may fail on hard programming tasks. The user experiences are good with overall average number of attempts being 1.61 and the average time of completion being 47.02 seconds. Our experiments also found that the weakest aspect of usability is conciseness, which has a score of 3.80 out of 5. Our experiment also shows that it is hard for human developers to learn from experiences to improve the skill of using ChatGPT to generate code.
@@ I Introduction
The past few years have seen a rapid growth of machine learning (ML) techniques in the development of large language models (LLMs). The most advanced LLMs like ChatGPT [^1] https://openai.com/chatgpt and Gemini [^2] https://deepmind.google/technologies/gemini have demonstrated their impressive performance in completing natural language processing (NLP) tasks [1]. Although their primary goal is for NLP, they have also shown the capable of performing various programming tasks with input in natural language and/or partially completed program code, because some of them have included program codes in their training dataset in addition to large volumes of natural language texts. Typical examples of such LLMs include OpenAI’s GPT-3.5 underlying ChatGPT and Google’s LaMDA model underlying Bard. Another approach to achieve the capability of performing coding tasks is through fine tuning of a pretrained LLM with program code datasets to build special purpose intelligent programming tools. Well-known examples of such systems include the Codex model underlying GitHub’s Copilot. Many ML models have also been purposely built for various programming tasks. For example, AlphaCode [2], InCoder [3], StarCoder [4], PolyCoder [5], Replit’s Ghostwrite [^3] https://replit.com/ and Tabnine [^4] https://www.tabnine.com/ have been developed for code completion, code generation, code comment summary generation, and debugging (code correction) tasks. ChatUNitTest [6], TestPilot [7], etc. are for the generation of test codes; see [8] for a survey. It is widely believed that it is becoming practical to automatically generate program code from natural language inputs; see, for example, [9, 10, 11].
@@
The paper is organized as follows. Section II reviews related works and discusses the problems associated with the existing methods of testing and evaluation of LLMs’ capability of coding. Section III presents the proposed user centric methodology. Section IV is devoted to the experiment results and the main findings. Section V concludes the paper with a summary, an analysis of the potential threats to the validity and generalizability of the experiment results, and a discussion of the directions for future work.
@@ II Related Work
II-A 2.1 General Evaluation of LLMs
@@
For example, the recent launch of Gemini [1] reports the performances of two versions of the ML model, i.e. Gemini-Ultra and Gemini-Pro, on many tasks of multi-modal language processing, which include multi-task language understanding, general reasoning, mathematical reasoning, and program code generation. It is compared with many front runners of LLMs, such as GPT-4, GPT-3.5, PalM 2-L, Inflection-2, Grok 1 and LLAMA-2. Table I lists some commonly used text datasets to benchmark LLMs on natural language text processing tasks. They are also used by Gemini team [1].
@@
MMLU [15]
Multichoice questions in 57 professional and academic subjects.
GSM8K[16]
Graduate school math questions.
MATH [17]
Mathematics problems across 5 difficulty levels and 7 subdisciplines.
BIG-Bench-Hard[18]
Hard Big-Bench tasks written as CoT problems.
HumanEval[19]
Python coding tasks.
Natural2Code [1]
Python code generation problems.
DROP[20]
Reading comprehension and arithmetic problems.
HellaSwag[21]
Common sense multiple-choice questions.
WMT23[22]
Machine translation problems.
@@ II-B Evaluation of LLMs on Code Generation
As discussed in the previous subsection, the capability of code generation is included in many of the general evaluation of LLMs. Much more work on the evaluation of LLMs’ capability of code generation has also been reported in the literature with the development of intelligent coding tools. Several datasets have been constructed and corresponding evaluation methods advanced.
The APPS dataset developed by Hendrycks et al. in 2021 [24] is perhaps the first benchmark for code generation from natural language input as functional specification. Earlier works have been reported in the literature for other coding tasks. For example, in year 2019, Kulal et al. evaluated the SPoC model, which generates code from pseudocode written in natural language [25]. In year 2020, Lachaux et al. evaluated a ML model for code translation [26].
@@
How to check the correctness of generated code automatically is an important issue. In the evaluations of NLP capabilities, the common approach is matching output with reference solutions either exactly or fuzzily, for example, using the ROUGE and BLEU metrics. However, for code generation, ROUGE and BLEU cannot capture the semantics of code. For example, Hendrycks et al. compared BLEU score with testing correctness [24]. They found that “BLEU increases as problem sources become more difficult, even though models actually perform worse on harder problems. Moreover, worse models can have similar or higher BLEU scores[24]. Chen et al. also pointed out that similarity metrics like BLEU score may not be a reliable indicator of functional correctness [19]. They also provided evidence that functionally inequivalent codes often have high BLEU scores. Ren et al. revised the BLEU metric and proposed CodeBLEU [27] to address the problem that the BLEU metric does not take the features of programming language syntax into consideration. However, the heart of the problem is that a coding problem can have many functionally equivalent solutions that are syntactically dissimilar to each other. In [31], Lai et al. used a much-relaxed form of similarity metric called surface-form constraints as a supplement to test correctness criterion discussed below. A surface-form constraint that requires the presence/absence of certain specific APIs and/or the keywords to occur in the solution code.
@@
An alternative to the uses of similarity metrics is to test the correctness of generated code. Borrowing the ideas of test-driving software development methodology, Kulal et al. proposed “pass test cases” as the correctness criterion to emphasize the importance of functional correctness over text similarity. Each code generation problem is therefore associated with a set of test cases. Du et al. further requires that the set of test cases to be highly adequate and uses test coverages of the reference solutions as measures of the quality of the dataset [28]. In [29], Liu et al. pointed out that the existing benchmark and the approach to evaluating code generation capability suffers from two problems. First, the test cases included in the task as the standard of correctness are insufficient to be a reliable criterion of correctness. Second, the natural language specifications of the tasks are often not precise and complete enough for LLM to process. Their solution to the problems is to generate test cases automatically and check the correctness of LLM output code against the output from the ground-truth solutions.
@@
The second milestone that Kuala et al. set is now known as the $pass@k$ metric. Kulal et al. asked the SPoC model to produce 100 solutions and considered the task of code generation is successfully completed if one of these solutions is correct in the terms of passing all test cases. This led to the $pass@k$ performance metrics, i.e., probability of completing tasks in $k$ trials. However, $pass@k$ tends to produce large variance [19]. The problem is addressed by Chen et al. in the evaluation of Codex [19]. They asked the model to produce a large number $n$ of solutions on each problem, where $n=200$, and counted the number $c$ of solutions that pass the correctness test. They used $n$ and $c$ to calculate an unbiased estimator of the $pass@k$ metric for $k<n$.
@@
Hendrycks et al. used APPS to evaluate several GPT models. They used two performance metrics: test case coverage and strict accuracy. The former is the average proportions of test cases on which the model passes the test. The latter is the ratio of the problems that the model passes all test cases. They calculated the overall performance of GPT models using $pass@1$ and $pass@5$.
@@
Du et al.’s ClassEval dataset [28] is a benchmark for generating code at class level while APPS and HumanEval are all benchmarks for code generation at standalone function or method level. Similar to APPS and HumanEval, each coding task in ClassEval comprises three elements: a description for the target class, a set of test cases for verifying the correctness of the generated code, and a canonical solution that acts as a reference solution. The description of the target class is a class skeleton with signatures of the methods of the class annotated with natural language texts to specify the functional requirements of the target class and methods and to describe the meanings of the parameters. The ClassEval dataset contains 100 problems to generate 100 classes with 412 methods. The data are manually curated from real-world projects and existing data from HumanEval and MBPP. An important feature of ClassEval is that they emphasized the adequacy of the test cases for the problem to facilitate reliable correctness checking of the generated code. On average, ClassEval’s test suites are of $98.2\%$ branch coverage and $99.75$ statement coverage.
@@
Lai et al.’s DS-1000 is a code generation benchmark specialized in Python code for data science problems [31]. It contains 1000 coding problems collected from Stack Overflow to reflect natural or real-world coding problems. The problems are also manually modified to create executable context, add reference solutions and additional test cases, and rewrite LLM unreadable data such as figures, etc. Their evaluation of LMMs employed test correctness as well as surface-form constraints that require the presence/absence of certain specific APIs as the keywords.
@@ II-C Analysis of Existing Work
A common feature of existing work on the evaluation of LLMs is that one figure is used to represent a LLM’s performance on one type of tasks such as mathematical reasoning, text comprehension, and code generation, etc. The figure is calculated from testing the ML model on a benchmark dataset using one metric. The most important advantage of the approach is that it offers a means of relatively objective comparison of different ML techniques and models. Such benchmarking has played a significant role in the research on ML techniques as pointed out by Martínez-Plumed et al. [32].
@@
However, the evaluations of LLMs have not reflected their usability from users’ point of view. For example, ChatGPT is popular to many users while its evaluation scores seem poor [33]. Codex only achieved a score of $28.8\%$ on $pass@100$ [19] while many users have highly appraised its coding capability. While human testers are in favour of ChatGPT’s output on text summary tasks over other state of art ML models, its evaluation result figure based on ROUGE metric shows the other way [13]. For the following reasons, such evaluation results may not reflect the usability of the models in real usage scenarios.
@@
First, to be a true reflection of the performance of a ML model in a specific real usage, the dataset used in the evaluation must have the same distribution as the users’ input profile. From the constructions of existing datasets, it is hard to see how these datasets represent user input profiles.
@@
Second, the users of a ML model may have different usage scenarios. The input profiles in different use scenarios have different distributions. For example, a senior software developer may have more complicated code generation tasks while junior programmers usually have less difficult coding tasks. A front-end developer may be more likely to produce GUI program code, while a back-end developer would have more tasks to write database related code, and a data analyst would need to generate more statistics code. Therefore, their input profile will have different distributions. Consequently, one ML model having a high performance for one user may perform poorly from another user’s point of view if their usages are very different. Even if a dataset that well represents the overall usage of the LLM for the average distribution of input profiles by all users, the figure obtained from such a benchmark will not reflect the usability for a specific type of users. It is desirable to know the performances of the LLM for different types of users in different scenarios. Moreover, for many natural language processing tasks, such as to write poems and to engage in dialogues with a human user, there is no objective correctness criterion. A solution that is good enough for one user could be far from satisfactory from another user’s point of view. The evaluation of ML models on such types of tasks is inevitably subjective. Using one decimal number to represent the performance of a ML model can hardly represent all users’ assessment.
@@
The above problems also occur in the evaluations of LLM’s capability of code generation. However, there are also a few issues specific to code generation.
@@
First, the quality attributes evaluated in the existing work are mostly focused on correctness of the code. There is a wide range of other important quality attributes because code generation is only a part of the software engineering process. Other quality attributes, like readability, well structuredness, logic clarity, etc., are important for software maintenance, evolution and reuse. They should also be assessed. Such quality attributes are also particularly important for intelligent interactive coding environments based on reinforcement learning such as Yang et al.’s recent work on InterCode [34].
@@
Second, a piece of code generated by a LLM is often classified in a binary way, i.e., the code is either correct or incorrect. In the real uses of LLMs for code generation, developers mostly use the generated codes as bases for further manual revision [9, 10, 11]. An incorrect code can also be useful if it can be easily revised into a good solution to the problem. It is desirable to evaluate the performance taking into consideration how close are the generated codes to the required results and how easy to make such changes to get a satisfactory solution.
@@
Moreover, the evaluation is a process of “one attempt per test case”. That is, on each test case, the LLM is invoked with one input, no matter whether it is a one-shot or multiple shot query [35], or multi-solution/multi-chances query (i.e. to generate many solutions on the same input) and the $pass@k$ metric is used for $k\geq 1$. However, in the real uses of LLMs as a code generation tool, the developers make many attempts to modify their inputs until they are satisfied with the output or give up after many attempts. It is desirable to evaluate LLMs’ performance on how many such attempts one could expect until a good solution is generated.
Finally, in the context of using ChatGPT as a code generation tool, natural language input is the human-computer interface. It is widely recognized as a great feature of ChatGPT over other means of invoking LLMs for completing coding tasks. For example, for code completion, incomplete code is the input. For denugging or code correction, incorrect code and error messages are the input. For code comment and summary generation, the code is the input. However, formulating a natural language input has been proven a non-trivial task. This raises an open question on the learnability of using LLM models. That is, how easy will the users learn formulating effective natural language inputs through the experiences of using a LLM like ChatGPT? Such evaluation is missing in the existing work on evaluation of LLMs.
III Methodology

### v3

# User Centric Evaluation of Code Generation Tools

@@ Abstract
With the rapid advance of machine learning (ML) technology, large language models (LLMs) are increasingly explored as an intelligent tool to generate program code from natural language specifications. However, existing evaluations of LLMs have focused on their capabilities in comparison with humans. It is desirable to evaluate their usability when deciding on whether to use a LLM in software production. This paper proposes a user centric method for this purpose. It includes metadata in the test cases of a benchmark to describe their usages, conducts testing in a multi-attempt process that mimics the uses of LLMs, measures LLM generated solutions on a set of quality attributes that reflect usability, and evaluates the performance based on user experiences in the uses of LLMs as a tool.
The paper also reports a case study with the method in the evaluation of ChatGPT’s usability as a code generation tool for the R programming language. Our experiments demonstrated that ChatGPT is highly useful for generating R program code although it may fail on hard programming tasks. The user experiences are good with overall average number of attempts being 1.61 and the average time of completion being 47.02 seconds. Our experiments also found that the weakest aspect of usability is conciseness, which has a score of 3.80 out of 5.
@@ I Introduction
The past few years have seen a rapid growth of machine learning (ML) techniques in the development of large language models (LLMs). The most advanced LLMs like ChatGPT [^1] https://openai.com/chatgpt and Gemini [^2] https://deepmind.google/technologies/gemini have demonstrated their impressive performance in completing natural language processing (NLP) tasks [1]. Although their primary goal is for NLP, they are also capable of performing various programming tasks with input in natural language and/or partially completed program code, because some of them have included program codes in their training dataset in addition to large volumes of natural language texts. Typical examples of such LLMs include OpenAI’s GPT-3.5 underlying ChatGPT and Google’s LaMDA model underlying Bard. Another approach to achieve the capability of performing coding tasks is through fine tuning of a pretrained LLM with program code datasets to build special purpose intelligent programming tools. Well-known examples of such systems include the Codex model underlying GitHub’s Copilot. Many ML models have also been purposely built for various programming tasks. For example, AlphaCode [2], InCoder [3], StarCoder [4], PolyCoder [5], Replit’s Ghostwrite [^3] https://replit.com/ and Tabnine [^4] https://www.tabnine.com/ have been developed for code completion, code generation, code comment summary generation, and debugging (code correction) tasks. ChatUNitTest [6], TestPilot [7], etc. are for the generation of test codes; see [8] for a survey. It is widely believed that it is becoming practical to automatically generate program code from natural language inputs; see, for example, [9, 10, 11].
@@
The paper is organized as follows. Section II reviews related works and discusses the problems associated with the existing methods of testing and evaluation of LLMs’ capability of coding. Section III presents the proposed user centric methodology. Section IV and V report a case study with ChatGPT to generate R program code. Section VI compares the proposed method with the existing works. Section VII concludes the paper with a summary and a discussion of the directions for future work.
@@ II Related Work
II-A General Evaluation of LLMs
@@
For example, the recent launch of Gemini [1] reports the performances of two versions of the ML model, i.e. Gemini-Ultra and Gemini-Pro, on many tasks of multi-modal language processing, which include multi-task language understanding, general reasoning, mathematical reasoning, and program code generation. It is compared with many front runners of LLMs, such as GPT-4, GPT-3.5, PalM 2-L, Inflection-2, Grok 1 and LLAMA-2. Table I lists some commonly used text datasets to benchmark LLMs on natural language text processing tasks. They are also used by the Gemini team [1].
@@
MMLU [15] Multichoice questions in 57 professional and academic subjects.
GSM8K[16] Graduate school math questions.
MATH [17] Mathematics problems across 5 difficulty levels and 7 subdisciplines.
BIG-Bench-Hard[18] Hard Big-Bench tasks written as CoT problems.
HumanEval[19] Python coding tasks.
Natural2Code [1] Python code generation problems.
DROP[20] Reading comprehension and arithmetic problems.
HellaSwag[21] Common sense multiple-choice questions.
WMT23[22] Machine translation problems.
@@ II-B Evaluation of LLMs on Code Generation
As discussed in the previous subsection, the capability of code generation is included in many of the general evaluations of LLMs. Much more work on the evaluation of LLMs’ capability of code generation has also been reported in the literature with the development of intelligent coding tools. Several datasets have been constructed and corresponding evaluation methods advanced.
The APPS dataset developed by Hendrycks et al. in 2021 [24] is perhaps the first benchmark specifically constructed for evaluating code generation from natural language input as functional specification. Earlier works have been reported in the literature for other coding tasks. For example, in 2019, Kulal et al. evaluated the SPoC model, which generates code from pseudocode written in natural language [25]. In 2020, Lachaux et al. evaluated a ML model for code translation [26].
@@
How to check the correctness of generated code automatically is an important issue. In the evaluations of NLP capabilities, the common approach is matching output with reference solutions either exactly or fuzzily, for example, using the ROUGE and BLEU metrics. However, for code generation, ROUGE and BLEU cannot capture the semantics of code. For example, Hendrycks et al. compared BLEU score with testing correctness [24]. They found that “BLEU increases as problem sources become more difficult, even though models actually perform worse on harder problems. Moreover, worse models can have similar or higher BLEU scores[24]. Chen et al. also pointed out that similarity metrics like BLEU score may not be a reliable indicator of functional correctness [19]. They also provided evidence that functionally inequivalent codes often have high BLEU scores. Ren et al. revised the BLEU metric and proposed CodeBLEU [27] to address the problem that the BLEU metric does not take the features of programming language syntax into consideration. However, the heart of the problem is that a coding problem can have many functionally equivalent solutions that are syntactically dissimilar to each other. In [31], as a supplement to test correctness criterion discussed below, Lai et al. used a much-relaxed form of similarity metric called surface-form constraints, where each constraint requires the presence/absence of certain specific APIs and/or the keywords to occur in the solution code.
@@
An alternative to the use of similarity metrics is to test the correctness of generated code. Borrowing the ideas of test-driving software development methodology, Kulal et al. proposed “pass test cases” as the correctness criterion to emphasize the importance of functional correctness over text similarity. Each code generation problem is therefore associated with a set of test cases. Du et al. further requires that the set of test cases to be highly adequate and uses test coverages of the reference solutions as measures of the quality of the dataset [28]. In [29], Liu et al. pointed out that the existing benchmark and the approach to evaluating code generation capability suffers from two problems. First, the test cases included in the task as the standard of correctness are insufficient to be a reliable criterion of correctness. Second, the natural language specifications of the tasks are often not precise and complete enough for LLM to process. Their solution to the problems is to generate test cases automatically and check the correctness of LLM output code against the output from the ground-truth solutions.
@@
The second milestone that Kuala et al. set is now known as the $pass@k$ metric. Kulal et al. asked the SPoC model to produce 100 solutions and considered the task of code generation is successfully completed if one of these solutions is correct in the terms of passing all test cases. This led to the $pass@k$ performance metrics, i.e., probability of completing tasks in $k$ trials. However, a straightforward calculation of $pass@k$ tends to produce large variance [19]. This problem is addressed by Chen et al. in the evaluation of Codex [19]. They asked the model to produce a large number $n$ of solutions on each problem, where $n=200$, and counted the number $c$ of solutions that pass the correctness test. They devised a formula to calculate an unbiased estimator of the $pass@k$ metric from the values of $n$ and $c$ for $k<n$.
@@
Hendrycks et al. used APPS to evaluate several GPT models. They used two performance metrics: test case coverage and strict accuracy. The former is the average proportions of test cases on which the generated solution passes the test. Therefore, it is better to be called average pass rate. The latter is the ratio of the problems that the generated solution passes all test cases. They calculated the overall performance of GPT models using $pass@1$ and $pass@5$.
@@
Du et al.’s ClassEval dataset [28] is a benchmark for generating code at class level while APPS and HumanEval are benchmarks for code generation at standalone function or method level. Similar to APPS and HumanEval, each coding task in ClassEval comprises three elements: a description for the target class, a set of test cases for verifying the correctness of the generated code, and a canonical solution that acts as a reference solution. The description of the target class is a class skeleton with signatures of the methods of the class annotated with natural language texts to specify the functional requirements of the target class and methods and to describe the meanings of the parameters. The ClassEval dataset contains 100 problems to generate 100 classes with 412 methods. The data are manually curated from real-world projects and existing data from HumanEval and MBPP. An important feature of ClassEval is that they emphasized the adequacy of the test cases for the problem to facilitate reliable correctness checking of the generated code. On average, ClassEval’s test suites are of $98.2\%$ branch coverage and $99.75\%$ statement coverage.
@@
Lai et al.’s DS-1000 is a code generation benchmark specialized for data science problems in Python[31]. It contains 1000 coding problems collected from Stack Overflow to reflect natural or real-world coding problems. The problems are also manually modified to create executable context, add reference solutions and additional test cases, and rewrite LLM unreadable data such as figures, etc. Their evaluation of LMMs employed test correctness as well as surface-form constraints that require the presence/absence of certain specific APIs as the keywords.
@@ II-C Analysis of Existing Work
A common feature of existing work on the evaluation of LLMs is that a single score is used to represent a LLM’s performance on one type of tasks such as mathematical reasoning, text comprehension, and code generation, etc. The score is calculated from testing the ML model on a benchmark dataset using one metric. The most important advantage of the approach is that it offers a means of relatively objective comparison of different ML techniques and models, and sometimes with human beings. Such benchmarking has played a significant role in the research on ML techniques as pointed out by Martínez-Plumed et al.[32].
@@
However, the evaluations of LLMs have not reflected their usability as intelligent tools from the users’ point of view. For example, ChatGPT is popular to many users while its evaluation scores seem poor [33]. Codex only achieved a score of $28.8\%$ on $pass@100$[19] while many users have highly appraised its coding capability. While human testers are in favour of ChatGPT’s output on text summary tasks over other state of art ML models, its evaluation score based on ROUGE metric shows the other way[13]. For the following reasons, such evaluation results may not reflect the usability of the models in real usage scenarios.
@@
First, to be a fair reflection of the performance of a ML model in a specific usage, the dataset used in the evaluation must have the same distribution as the users’ input profile. From the constructions of existing datasets, it is hard to see how these datasets represent user input profiles.
@@
Second, the users of a ML model may have different use scenarios. The input profiles in different use scenarios have different distributions. For example, a senior software developer may have more complicated code generation tasks while junior programmers usually have less difficult coding tasks. A front-end developer may be more likely to produce GUI program code, while a back-end developer would have more tasks to write database related code, and a data analyst would need to generate more statistical data processing code. Therefore, their input profile will have different distributions. Consequently, one ML model having a high performance for one user may perform poorly from another user’s point of view if their use scenarios are very different. Even if a dataset that well represents the overall usage of the LLM for the average distribution of input profiles by all users, the figure obtained from such a benchmark will not reflect the usability for a specific type of users. It is desirable to know the performances of the LLM for different types of users in different scenarios.
Moreover, for many natural language processing tasks, such as to write poems and to engage in dialogues with a human user, there is no objective correctness criterion. A solution that is good enough for one user could be far from satisfactory from another user’s point of view. The evaluation of ML models on such types of tasks is inevitably subjective. Using one score to represent the performance of a ML model can hardly represent all users’ assessment.
@@
The above problems also occur in the evaluations of a LLM’s capability of code generation. However, there are also a few issues specific to code generation.
@@
First, the quality attributes evaluated in the existing work are mostly focused on functional correctness of the code. There is a wide range of other important quality attributes because code generation is only a part of the software engineering process. Other quality attributes, like readability, well structuredness, logic clarity, etc., are important for software maintenance, evolution, and reuse. They should also be assessed. Such quality attributes are also particularly important for intelligent interactive coding environments based on reinforcement learning such as Yang et al.’s recent work on InterCode [34].
@@
Second, a piece of code generated by a LLM is often classified in a binary way, i.e., the code is either correct or incorrect. In the real uses of LLMs for code generation, developers mostly use the generated codes as bases for further manual revision [9, 10, 11]. An incorrect code can also be useful if it can be easily revised into a good solution to the problem. It is desirable to evaluate the performance taking into consideration how close are the generated codes to the required results and how easy to revise the generated code to make a good solution.
@@
Moreover, the testing and evaluation process can be characterised by “one attempt per problem”. That is, on each coding problem, the LLM is invoked with one input to generate one or many solutions, no matter whether the input is a one-shot or multiple shot query [35]. This process is reflected in the $pass@k$ metric. Even if $k>1$, it is still “one attempt” in the sense that it queries the model with only one input, although it tries to get multiple solutions, or equivalently, gives the model multiple chances to answer the same query. However, in the real uses of LLMs as a code generation tool, the developers make many attempts by modifying their inputs until they are satisfied with the output, or they give up after too many attempts. It is desirable to evaluate LLMs’ performance on how many such attempts one could expect until a solution good enough is generated.
III The Proposed Methodology